OpenStudy (anonymous):

The book defines the completes solution of Ax=b as X(particular) + X(nullspace). Can anyone elaborate further?

OpenStudy (anonymous):

x(particular) solves the given equation: Ax(particular) = b. And x(nullspace) satisfies: Ax(nullspace) = 0. So Ax(particular) + Ax(nullspace) = Ax(particular) + 0 = b. Hows that?

OpenStudy (anonymous):

By definition, any vectors in the nullspace N(A), satisfy Ax = 0. For example, say n was in the nullspace then An =0. Now, if you go back to the problem where you are looking for a solution x to the problem Ax = b and lets say you found one, call it p. In other words Ap = b. What the statement says it that using that solution we can find another one by using the vector in the nullspace. For example what about the vector p + n? A(p + n) = Ap + An = b + 0 = 0 We could also have used 10n or -3n or any other scalar value. When it talks about the "complete" solution, it means what are all of the possible solutions to Ax = b. This turns out to be any vector that is the sum of a particular solution and any combination of vectors in the nullspace N(A) since they all go factor to zero. In other words, if you find any particular solution, and you find a basis for the nullspace, then you've found all possible solutions. Dunno if that helps.